OpenStudy (anonymous):

Please help. The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank: 1 over 3π (x - 7)2 (x - 3) What does the factor 1 over 3(x - 7) represent?

OpenStudy (anonymous):

The total surface area of the tank One third of the radius of the tank One third of the height of the tank The total area of the base of the tank

OpenStudy (anonymous):

hey can you help me with this one @rosesmarie its the last question

OpenStudy (anonymous):

@OregonDuck @Peaches15

OpenStudy (anonymous):

@peachpi

OpenStudy (anonymous):

hello?

OpenStudy (rhr12):

take a look at it

OpenStudy (anonymous):

i should have asked, if \((x-3)\) is the height of tank, then \((x-7)\) is what? height = x-3 radius = h-4 = x-3-4 = x - 7 that means, (x-7) is the radius of tank so \(\frac{1}{3}(x-7)\) is One third the radius of tank